Suppose that you have been offered two packages. Package A contains 1000 call options at strike price of $182. Package B contains 100 call options at strike price of $80. Suppose stock price next year could be $100 with probability 0.75 and $200 with probability 0.25. Suppose that you are risk neutral. Which package would you choose
Answer #1


Payoff of 1 call option 1 year later in Package A=max(100-182,0)*0.75+max(200-100,0)*0.25=25
Total payoff from Package A=25*1000=25000
Payoff of 1 call option 1 year later in Package B=max(100-80,0)*0.75+max(200-80,0)*0.25=20
Total payoff from Package B=20*100=2000
If both package are offereed at the same price, we should accept Package A